# Uncertainty-Aware Surrogate-based Amortized Bayesian Inference for Computationally Expensive Models

### v1

@@ 4.2.3 Validation of Parameter Posterior
We validate the correctness of the surrogate-based ABI approaches and benchmark their performance against a reference solution. To this end, we compare the two surrogate-based ABI methods, SABI and UA-SABI, against the corresponding surrogate-based MCMC methods, Point and E-Post. We plot the posterior medians (circles) and their median deviations (lines) for the ABI methods along with the corresponding MCMC results, as well as the ECDF differences for all the methods in Fig. 4.
((a)) Recovery plots.
((b)) ECDF difference plots.
Figure 4: Recovery plots (left) and ECDF difference plots (right) plots comparing ABI methods to corresponding MCMC methods over 200 ground truth samples. In the ECDF difference plots, empirical ranks are shown in blue, $95\%$ confidence bands assuming calibration are shown in grey.
@@ 4.2.4 Runtime Comparison
In Fig. 6(a), we show the runtime of UA-SABI and E-Post for $\{5,10,15,20\}$ inference runs. We observe that UA-SABI’s runtime is nearly constant for the number of inference runs, while E-Post scales linearly. Based on this comparison, UA-SABI is already justified after around 9 inference runs.
@@ 4.3 Case Study 2: Carbon Dioxide ($\text{CO}_{2}$) Storage Model
In the second case study, we test our method on a real-world problem. Specifically, we consider a $\text{CO}_{2}$ storage benchmark [19]. In this test case, a non-linear hyperbolic partial differential equation models the two-phase flow of $\text{CO}_{2}$ in brine. It describes $\text{CO}_{2}$ injection, plume migration, pressure build-up, and the influence of uncertain porous medium properties in a deep saline aquifer. Given the $\text{CO}_{2}$ saturation as measurements, our aim is to infer 3 parameters of interest: injection rate of $\text{CO}_{2}$ (IR), relative permeability degree in the fractional flux function (PM), and porosity (PR) of the formation. Five separate Bayesian PCE models with sparsity-inducing priors [2] are trained for five different instants at days $\{20,40,60,80,100\}$ and at a fixed location $30$m from the injection well. A similar setup was studied in [26]. To avoid redundancy, we show results only for porosity.
@@ 4.3.1 Setup
In general, the setup for the $\text{CO}_{2}$ storage model follows the same structure as that described in Section 4.2.1. The priors of the input parameters are given as $\text{IR}\sim 6.4\times 10^{-4}\times(1+\text{Beta}(4,2))$, $\text{PM}\sim 2\times\text{Beta}(1.25,1.25)+2$, and $\text{PR}\sim\text{Beta}(2.4,9)$ [19]. For PCE training, we considered 64 Sobol sequence evaluations scaled to the input parameter prior ranges and constructed the polynomial basis with arbitrary polynomial chaos (aPC) based on the priors [25, 2].
In contrast to case study 1, we used an offline ABI training set, where the training data is pre-generated from $10^{4}$ parameter prior draws and fixed during the training process, allowing for direct comparison with a standard full-budget ABI trained on pre-generated simulation data. Further computational details are also given in Appendix C.
@@ 4.3.2 Validation of Parameter Posterior
We compare the inferred posterior samples obtained from a full-budget ABI ($N_{B}=10^{4}$), a low-budget ABI ($N_{B}=64$), SABI, and UA-SABI. Full-budget ABI was feasible due to the already available data [18, 26]. Therefore, recovery plots were chosen to validate the quality of our results relative to standard ABI. The porosity results are shown in Fig. 5.
Recovery plots in Fig. 5(a) again show that low-budget ABI struggles to recover the ground truth parameter, particularly near domain boundaries. SABI also produces poor estimates and, moreover, fails to capture uncertainty in the inferred parameter posteriors. This overconfidence results in miscalibrated posteriors, confirmed by the two corresponding ECDF difference plots in Fig. 5(b).
In contrast, UA-SABI performs comparably to full-budget ABI while accounting for the additional uncertainty introduced by using a surrogate to generate training data. According to the ECDF difference plots, it produces a well-calibrated posterior for porosity. Recovery plots between ABI and MCMC for surrogate-based methods and corresponding ECDF difference plots are provided in Appendix D.
((a)) Recovery plots of porosity.
((b)) ECDF difference plots of porosity. Only full-budget ABI and UA-SABI are well calibrated.
Figure 5: Recovery plots (top) and ECDF difference plots (bottom) for full-budget ABI, low-budget ABI, SABI, and UA-SABI over 200 ground truth samples. In the ECDF difference plots, empirical ranks are shown in blue, $95\%$ confidence bands assuming calibration are shown in grey.
@@ 4.3.3 Runtime Comparison
Also, for the $\text{CO}_{2}$ storage model, we compare the runtimes, following the same approach as in Section 4.2.4. The experiments were run on a computing cluster with two AMD EPYC 7551 CPUs (totaling 64 physical cores) to speed up E-Post through parallelization.
Figure 6(b) presents the measured runtimes of UA-SABI and E-Post for $\{5,6,7,8\}$ inference runs, whereby E-Post was parallelized on 16 cores. We observe a break-even point between 6 and 7 inference runs, indicating that UA-SABI becomes justified after 7 runs. Despite using more cores for E-Post, the training costs are amortized earlier for a more expensive model.
((a)) Runtimes for LogSin for $\{5,10,15,20\}$ runs.
((b)) Runtimes for $\text{CO}_{2}$ for $\{5,6,7,8\}$ runs.
@@ 5 Summary and Outlook
In this work, we introduced Uncertainty-Aware Surrogate-based Amortized Bayesian Inference (UA-SABI) – a framework designed to enable efficient and reliable ABI for computationally expensive models. UA-SABI combines surrogate modeling and ABI while explicitly quantifying and propagating surrogate uncertainties through the inference process. This addresses a core limitation of existing approaches: while surrogate models can reduce the cost of generating training data for ABI, ignoring their approximation error often leads to overconfident and misleading posteriors. By incorporating uncertainty awareness, UA-SABI enables better calibrated and reliable inference, even under tight computational constraints.
@@
We validated UA-SABI in both a simple toy example and a real-world problem of modeling $\text{CO}_{2}$ storage, highlighting its ability to produce well-calibrated posterior estimates that match those obtained with MCMC-based methods. Our experiments demonstrate the importance of sufficient training data, as low-budget ABI produces erroneous posteriors. Also, they show the importance of uncertainty propagation when using a surrogate: SABI results in overconfident posteriors, whereas UA-SABI correctly reflects model uncertainty in its predictions. Moreover, we showed that the upfront computational effort of training UA-SABI is quickly offset in scenarios involving repeated inference, with amortization becoming beneficial already after a few inference runs.
Overall, UA-SABI offers a robust and efficient solution for more reliable and scalable inference in computationally demanding settings. Future work could include testing UA-SABI on models with higher-dimensional parameter spaces to further evaluate its scalability and performance.
Acknowledgments and Disclosure of Funding
@@
We thank the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) for supporting this work by funding – EXC2075 – 390740016 under Germany’s Excellence Strategy and the Collaborative Research Centre SFB 1313, Project Number 327154368. We acknowledge the support by the Stuttgart Center for Simulation Science (SimTech). We further acknowledge the support of the DFG Collaborative Research Center 391 (Spatio-Temporal Statistics for the Transition of Energy and Transport) – 520388526. Additionally, we thank Dr. Ilja Kröker for his help and for providing the data.
@@ Appendix A Pseudocode
2:Simulation model $M(\mathbf{x},\boldsymbol{\omega})$, prior $p(\mathbf{c},\sigma)$, likelihood $p(\mathbf{y}\mid\mathbf{x},\boldsymbol{\omega},\mathbf{c},\sigma)$
@@
4:Choose inputs and parameters $\{(\mathbf{x}^{(i)},\boldsymbol{\omega}^{(i)})\}_{i=1}^{N_{T}}$
@@
6:     Evaluate simulation model $\mathbf{y}^{(i)}=M(\mathbf{x}^{(i)},\boldsymbol{\omega}^{(i)})$
@@
8:Construct training dataset $D_{T}=\{(\mathbf{x}^{(i)},\boldsymbol{\omega}^{(i)},\mathbf{y}^{(i)})\}_{i=1}^{N_{T}}$
@@
$p(\mathbf{c},\sigma\mid D_{T})\propto\prod_{i=1}^{N_{T}}p(\mathbf{y}^{(i)}\mid\mathbf{x}^{(i)},\boldsymbol{\omega}^{(i)},\mathbf{c},\sigma)\,p(\mathbf{c},\sigma)$
@@
11:Prior $p(\mathbf{x},\boldsymbol{\omega})$, posterior $p(\mathbf{c},\sigma\mid D_{T})$
12:Trained summary and inference network parameters $\widehat{\boldsymbol{\theta}},\widehat{\boldsymbol{\varphi}}$
@@
15:         Sample inputs and parameters from prior $(\mathbf{x}^{(i)},\boldsymbol{\omega}^{(i)})\sim p(\mathbf{x},\boldsymbol{\omega})$
@@
17:         Evaluate surrogate $\widetilde{\mathbf{y}}^{(i)}=\widetilde{M}_{\mathbf{c}^{(i)}}(\mathbf{x}^{(i)},\boldsymbol{\omega}^{(i)})$
@@
19:         Pass $(\mathbf{x}^{(i)},\widetilde{\mathbf{y}}_{\epsilon}^{(i)})$ through summary network $\mathbf{s}^{(i)}=S_{\boldsymbol{\theta}}(\mathbf{x}^{(i)},\widetilde{\mathbf{y}}_{\epsilon}^{(i)})$
20:         Pass $(\boldsymbol{\omega}^{(i)},\mathbf{s}^{(i)})$ through inference network $q_{\boldsymbol{\varphi}}(\boldsymbol{\omega}^{(i)}\mid\mathbf{s}^{(i)})$
21:         Compute loss from Eq. 2 and update network parameters $\boldsymbol{\theta},\boldsymbol{\varphi}$

### v2

@@ 4.2.3 Validation of Parameter Posterior
We validate the correctness of the surrogate-based ABI approaches and benchmark their performance against a reference solution. To this end, we compare the two surrogate-based ABI methods, SABI and UA-SABI, against the corresponding surrogate-based MCMC methods, Point and E-Post. We plot the posterior medians (circles) and their median deviations (lines) for the ABI methods along with the corresponding MCMC results, as well as the ECDF differences for all the methods in Fig.˜4.
(a) Recovery plots.
(b) ECDF difference plots.
Figure 4: LogSin recovery plots (left) and ECDF difference plots (right) comparing ABI methods to corresponding MCMC methods over 200 ground truth samples. In the ECDF difference plots, empirical ranks are shown in blue, $95\%$ confidence bands assuming calibration are shown in grey.
@@ 4.2.4 Runtime Comparison
In Fig.˜6(a), we show the runtime of UA-SABI and E-Post for $\{5,10,15,20\}$ inference runs. We observe that UA-SABI’s runtime is nearly constant for the number of inference runs, while E-Post scales linearly. Based on this comparison, UA-SABI is already justified after around 9 inference runs.
@@ 4.3 Case Study 2: Carbon Dioxide ($\text{CO}_{2}$) Storage Model
In the second case study, we test our method on a virtual benchmark that represents a real-world problem. Specifically, we consider a $\text{CO}_{2}$ storage benchmark (Köppel et al., 2019). In this test case, a non-linear hyperbolic partial differential equation models the two-phase flow of $\text{CO}_{2}$ in brine. It describes $\text{CO}_{2}$ injection, plume migration, pressure build-up, and the influence of uncertain porous medium properties in a deep saline aquifer. Given the $\text{CO}_{2}$ saturation as measurements, our aim is to infer three parameters of interest: injection rate of $\text{CO}_{2}$ (IR), relative permeability degree in the fractional flux function (PM), and porosity (PR) of the formation. This test case allows us to analyze the performance of UA-SABI in inference for hydrosystem models. The underlying physical model exhibits strong discontinuities at certain combinations of parameter values and space/time coordinates, which inherently limit the approximation performance of PCE-based surrogate models. However, this scenario provides an opportunity to test UA-SABI under challenging conditions where the surrogate model is strongly misspecified. Five separate Bayesian PCE models with sparsity-inducing priors (Bürkner et al., 2023) are trained for five different instants at days $\{20,40,60,80,100\}$ and at a fixed location 3030m from the injection well. A similar setup was studied in (Oladyshkin et al., 2020). We show parameter posterior results for porosity, the most informative parameter of the $\text{CO}_{2}$ storage model.
@@ 4.3.1 Setup
In general, the setup for the $\text{CO}_{2}$ storage model follows the same structure as that described in Section˜4.2.1. The priors of the input parameters are given as $\text{IR}\sim 6.4\times 10^{-4}\times(1+\text{Beta}(4,2))$, $\text{PM}\sim 2\times\text{Beta}(1.25,1.25)+2$, and $\text{PR}\sim\text{Beta}(2.4,9)$ (Köppel et al., 2019). For PCE training, we considered 64 Sobol sequence evaluations scaled to the input parameter prior ranges and constructed the polynomial basis with arbitrary polynomial chaos (aPC) based on the priors (Oladyshkin & Nowak, 2012; Bürkner et al., 2023).
In contrast to case study 1, we used an offline ABI training set, where the training data is pre-generated from $10^{4}$ parameter prior draws and fixed during the training process, allowing for direct comparison with a standard full-budget ABI trained on pre-generated simulation data. Further computational details are given in Appendix˜C.
@@ 4.3.2 Validation of Parameter Posterior
We compare the inferred posterior samples obtained from a full-budget ABI ($N_{B}=10^{4}$), a low-budget ABI ($N_{B}=64$), SABI, and UA-SABI. Full-budget ABI was feasible due to the already available data (Köppel et al., 2017; Oladyshkin et al., 2020). Therefore, recovery plots were chosen to validate the quality of our results relative to standard ABI (Fig.˜5).
Recovery plots in Fig.˜5(a) again show that low-budget ABI struggles to recover the ground truth parameter, particularly near domain boundaries. SABI also produces poor estimates and, moreover, fails to capture uncertainty in the inferred parameter posteriors. This overconfidence results in miscalibrated posteriors, confirmed by the two corresponding ECDF difference plots in Fig.˜5(b).
In contrast, UA-SABI performs comparably to full-budget ABI while accounting for the additional uncertainty introduced by using a surrogate to generate training data. According to the ECDF difference plots, it produces a well-calibrated posterior for porosity. Additionally, recovery plots between ABI and MCMC for surrogate-based methods and ECDF difference plots for MCMC-based methods are provided in Appendix˜D. Despite the model’s discontinuities and the resulting surrogate misspecification, our method successfully quantifies and propagates surrogate uncertainty to the parameter posterior.
(a) Recovery plots of porosity.
(b) ECDF difference plots of porosity. Only full-budget ABI and UA-SABI are well calibrated.
Figure 5: $\text{CO}_{2}$ recovery plots (top) and ECDF difference plots (bottom) for full-budget ABI, low-budget ABI, SABI, and UA-SABI (from left to right) over 200 ground truth samples. In the ECDF difference plots, empirical ranks are shown in blue, $95\%$ confidence bands assuming calibration are shown in grey.
@@ 4.3.3 Runtime Comparison
Also, for the $\text{CO}_{2}$ storage model, we compare the runtimes, following the same approach as in Section˜4.2.4. The experiments were run on a computing cluster with two AMD EPYC 7551 CPUs (totaling 64 physical cores) to speed up E-Post through parallelization.
Figure˜6(b) presents the measured runtimes of UA-SABI and E-Post for $\{5,6,7,8\}$ inference runs, whereby E-Post was parallelized on 16 cores. We observe a break-even point between 6 and 7 inference runs, indicating that UA-SABI becomes justified after 7 runs. Despite using more cores for E-Post, the training costs are amortized earlier for a more expensive model.
(a) Runtimes for LogSin for $\{5,10,15,20\}$ runs.
(b) Runtimes for $\text{CO}_{2}$ for $\{5,6,7,8\}$ runs.
@@
4.4 Case Study 3: Microbially Induced Calcite Precipitation Model
In this case study, we transition from benchmarks to realistic conditions by using real measurement data. Case studies 1 and 2 introduced a toy example and a virtual benchmark, respectively. Case study 3 now focuses on posterior inference from observed measurements, where the primary goal is to evaluate the applicability of UA-SABI under real-world conditions. A runtime comparison for the MICP model on synthetic data is added in Section˜E.4.
The underlying model predicts microbially induced calcite precipitation (MICP). MICP is a reactive transport process, here taking place inside a porous medium. It includes at least biofilm, calcite, and the unreactive solid matrix as solid phases, water, and, in some cases, an additional fluid phase such as gas. Relevant in the aqueous phase are dissolved calcium, inorganic carbon, and urea. The complete set of relevant species is detailed in Hommel et al. (2015). The process is mediated by S. pasteurii, a bacterium that produces the enzyme urease. Urease catalyzes the hydrolysis of urea into ammonia and carbonic acid. In aqueous solution, ammonia consumes hydrogen ions ($H^{+}$), which raises the pH. The increase in pH alters carbonate equilibria: carbonic acid dissociates further, releasing additional $H^{+}$ and carbonate ions. This raises the concentration of dissolved carbonate species. When calcium ions are present, they combine with carbonate ions and trigger the precipitation of calcite within the pore space.
The experiment is performed on a $61\text{cm}$ sand-filled column with a $2.54\text{cm}$ diameter. Bacteria are first injected from the bottom, followed by an overnight no-flow period to allow biofilm formation. Biofilm growth is then stimulated with a $24\text{h}$ substrate injection. Afterwards, two pore volumes of a $0.33\text{mol}/\text{L}$ calcium–urea solution are injected at $10\text{ml}/\text{min}$, repeated every $24\text{h}$. Each injection is followed by a no-flow period for mineralization, and then a substrate injection to reactivate the biofilm (Hommel et al., 2015). This cycle is repeated 30 times. A more detailed description can be found in Hommel et al. (2015).
Measurements of the volume fraction of calcite on eight spatial points located at $3.81$, $11.43$, $19.05$, $26.67$, $34.29$, $41.91$, $49.53$, and $57.15\text{cm}$ distance from the bottom are available. After the last cycle, we build our surrogate based on four parameters of interest: the coefficient for preferential attachment to biomass ($c_{a,1},[\text{s}^{-1}]$), the coefficient for attachment to arbitrary surfaces ($c_{a,2},[\text{s}^{-1}]$), the dry mass density of biofilm ($\rho_{f},[\text{kg}/\text{m}^{3}]$), and the enzyme content of biomass ($k_{ub},[\text{kg}/\text{kg}]$).
The MICP case provides an opportunity to test UA-SABI on a model that is highly computationally expensive yet low-dimensional in parameter space, and for which real-world measurement data are available. We again employ PCE as the surrogate modeling technique, due to its demonstrated success for this model (Scheurer et al., 2021). In total, eight Bayesian PCEs are trained, one for each measurement location.
4.4.1 Setup
In general, the setup for the MICP model follows the same structure as that described in the former case studies. The priors of the input parameters are given as $c_{a,1}\sim\mathcal{U}(10^{-10},10^{-7})$, $c_{a,2}\sim\mathcal{U}(10^{-10},10^{-6})$, $\rho_{f}\sim\mathcal{U}(1,15)$, and $k_{ub}\sim\mathcal{U}(10^{-5},5\cdot 10^{-4})$ (Scheurer et al., 2021). For PCE training, 25 model evaluations within the prior ranges of input parameters are available, which are not necessarily optimal for surrogate training. We constructed the polynomial basis with aPC based on the available model evaluations (Oladyshkin & Nowak, 2012; Bürkner et al., 2023).
As in case study 2, we used an offline ABI training set of $10^{4}$ parameter prior draws. In contrast to case study 2, no pre-generated simulation data are available for these draws. Instead, output data can only be generated through the surrogate; thus, no comparison with a full-budget ABI is possible. Further computational details are given in Appendix˜C.
4.4.2 Validation of Surrogate Model and Parameter Inference Setup
Since inference will be performed using real measurement data, the first step is to evaluate how accurately the surrogate represents the underlying system. Figure˜7(a) presents the measurement data, the model evaluations used for surrogate training, and the surrogate predictions obtained from $10^{4}$ prior samples across the eight locations. The results indicate that the simulated model data—and consequently the surrogate—do not cover the region occupied by the measurement data. This makes the measurement data out-of-distribution (OOD) relative to the simulation model’s a priori predictions - a phenomenon that is suspected to occur frequently in dynamic systems modeling, although rarely investigated in detail.
(a) Surrogate output with original prior for $k_{ub}\sim\mathcal{U}(10^{-5},5\cdot 10^{-4})$.
(b) Surrogate output with adjusted prior for $k_{ub}\sim\mathcal{U}(10^{-5},5\cdot 10^{-3})$.
Figure 7: Measurement data for calcite, 25 model evaluations, and surrogate model output obtained with original prior (left) and adjusted prior (right) for $k_{ub}$ of the MICP model across eight locations. Surrogate predictions are obtained from $10^{4}$ prior samples; median and 90% confidence intervals are shown.
Neural networks are known to generalize poorly outside their training domain, resulting in poor posteriors obtained via ABI (Schmitt et al., 2024a), which is shown in Section˜E.1. Thus, we seek to transfer this challenge to the surrogate instead. To this end, we found that the surrogate priors used for parameter inference can be adjusted to ensure that the measurement data fall within the surrogate’s support. This approach leverages the generalization properties of the low-dimensional, regularized surrogate, avoiding dependence on the poor generalization abilities of neural networks, which are used within the ABI model during inference.
The results of a sensitivity analysis (more detailed in Section˜E.2) on the surrogate using Sobol indices (Sobol’, 1990) shows that the surrogate’s output at all locations is almost only sensitive to $k_{ub}$.
In Fig.˜8, the surrogate output is shown exemplarily at location index 3 by varying $k_{ub}$. It can be observed that the uncertainty of the surrogate increases when evaluating OOD inputs. Based on these results, we adapt the parameter inference prior range for $k_{ub}$, such that $k_{ub}\sim\mathcal{U}(10^{-5},5\cdot 10^{-3})$. This adjustment results in a prior predictive that covers the measurement data, as illustrated in Fig.˜7(b).
Figure 8: MICP surrogate output at location index 3 over $k_{ub}$. The output is obtained by evaluating the uncertainty-aware surrogate for $k_{ub}\in[10^{-5},5\cdot 10^{-3}]$ and keeping the other parameters fixed at their means. Median and 90% confidence intervals of the surrogate output are shown in green, the region of the original prior is shown in grey.
4.4.3 Comparison of Parameter Posterior
Since parameter inference is carried out on real measurement data, no ground truth is available. We therefore compare only the inferred posteriors obtained using low-budget ABI ($N_{B}=25$), SABI, UA-SABI, Point, and E-Post. For easy visual comparison, we applied kernel density estimation (KDE) with Scott’s bandwidth method to the posterior samples from each approach. Since $k_{ub}$ is the only sensitive parameter, we present its posterior estimates in Fig.˜9 to avoid redundancy, while posterior estimates for the remaining parameters are provided in Section˜E.3.
Figure 9: KDEs of 4,000 parameter posterior samples for $k_{ub}$ of the MICP model obtained with low-budget ABI, Point, E-Post, SABI, and UA-SABI.
Figure˜9 demonstrates that low-budget ABI, which is shown as a point estimate due to quasi-nonexistent variance in the posterior, exhibits a pronounced bias and substantially underestimates posterior uncertainty. This confirms again that training an ABI model with very limited data is not meaningful. Comparing the surrogate-based methods further highlights that explicitly quantifying surrogate uncertainty and propagating it through the inference process is reflected in the wider posterior estimates. To assess the validity of the obtained posteriors and the appropriateness of the quantified and propagated uncertainty, Fig.˜10 presents the posterior predictive distributions for calcite across the four surrogate-based methods.
(a) Point.
(b) SABI.
(c) E-Post.
(d) UA-SABI.
Figure 10: MICP posterior predictives for calcite for the four surrogate-based methods. Surrogate predictions are obtained from 4,000 posterior samples, median and 90% confidence intervals are shown.
Figure˜10(a) and Fig.˜10(b) clearly illustrate that not quantifying and propagating uncertainty results in overconfident posterior predictives. In contrast, as shown in Fig.˜10(c) and Fig.˜10(d), explicitly accounting for uncertainty produces confidence intervals that successfully cover the measurement data. Both methods yield comparable intervals, with UA-SABI producing a slightly broader interval as a higher posterior density in the out-of-distribution region of the most sensitive parameter $k_{ub}$ (Fig.˜9) yields increased surrogate predictive uncertainty.
Overall, this demonstrates that quantifying and propagating the uncertainty of a surrogate model enables ABI for highly computationally expensive models. Even when the available model simulations are suboptimal for surrogate training and far from the measurement data, accounting for uncertainty and promoting generalization in the surrogate model leads to more trustworthy results.
@@ 5 Summary and Outlook
In this work, we introduced Uncertainty-Aware Surrogate-based Amortized Bayesian Inference (UA-SABI) – a framework designed to enable efficient and reliable ABI for computationally expensive models. UA-SABI combines surrogate modeling and ABI while explicitly quantifying and propagating surrogate uncertainties through the inference process. This addresses a core limitation of existing approaches: while surrogate models can reduce the cost of generating training data for ABI, ignoring their approximation error often leads to overconfident and misleading posteriors. By incorporating uncertainty awareness, UA-SABI enables better-calibrated and reliable inference, even under tight computational constraints.
@@
We validated UA-SABI in both a simple toy example and two real-world problems of modeling $\text{CO}_{2}$ storage and MICP, highlighting its ability to produce well-calibrated posterior estimates that match those obtained with MCMC-based methods. Our experiments demonstrate the importance of sufficient training data, as low-budget ABI produces erroneous posteriors. Also, they show the importance of uncertainty propagation when using a surrogate: Surrogate-based ABI results in overconfident posteriors, whereas UA-SABI correctly reflects model uncertainty in its predictions. Moreover, we showed that the upfront computational effort of training UA-SABI is quickly offset in scenarios involving repeated inference, with amortization becoming beneficial already after a few inference runs. Additionally, the MICP experiment poses a challenging real-world scenario by having 1) only a limited number of model runs available, and 2) real measurement data that are not adequately captured by these model runs.
Overall, UA-SABI offers an efficient and reliable solution in settings that require repeated inference for computationally expensive models. In general, we conclude that using surrogates to generate training data can be an effective strategy for ABI when simulations are scarce or computationally expensive. Moreover, explicitly accounting for surrogate uncertainty in the training data improves ABI results, particularly by reducing overconfidence.
In this work, we applied UA-SABI only to low-dimensional, computationally expensive problems to keep results interpretable and avoid challenges that would distract from our scope - such challenges as they appear in surrogate training and parameter inference for high-dimensional models with only scarce available data. Future work should test the method on higher-dimensional parameter spaces to further evaluate its scalability and performance. Additionally, the surrogate is technically misspecified, as the simulator is not fully contained within the surrogate class. Thus, future work could integrate UA-SABI with recently proposed methods for detecting and mitigating model misspecification in ABI (Schmitt et al., 2024a; Dellaporta et al., 2022; Elsemüller et al., 2025), which are all naturally compatible with our framework.
Acknowledgments
@@
We thank the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) for supporting this work by funding – EXC2075 – 390740016 under Germany’s Excellence Strategy and the Collaborative Research Centre SFB 1313, Project Number 327154368. We acknowledge the support by the Stuttgart Center for Simulation Science (SimTech). We further acknowledge the support of the DFG Collaborative Research Center 391 (Spatio-Temporal Statistics for the Transition of Energy and Transport) – 520388526. Additionally, we thank Dr. Ilja Kröker for his help and and to both Dr. Kröker and Dr. Johannes Hommel for providing data.
@@ Appendix A Pseudocode
2:Simulation model $M(\mathbf{x},\bm{\omega})$, prior $p(\mathbf{c},\sigma)$, likelihood $p(\mathbf{y}\mid\mathbf{x},\bm{\omega},\mathbf{c},\sigma)$
@@
4:Choose inputs and parameters $\{(\mathbf{x}^{(i)},\bm{\omega}^{(i)})\}_{i=1}^{N_{T}}$
@@
6:  Evaluate simulation model $\mathbf{y}^{(i)}=M(\mathbf{x}^{(i)},\bm{\omega}^{(i)})$
@@
8:Construct training dataset $D_{T}=\{(\mathbf{x}^{(i)},\bm{\omega}^{(i)},\mathbf{y}^{(i)})\}_{i=1}^{N_{T}}$
@@
$p(\mathbf{c},\sigma\mid D_{T})\propto\prod_{i=1}^{N_{T}}p(\mathbf{y}^{(i)}\mid\mathbf{x}^{(i)},\bm{\omega}^{(i)},\mathbf{c},\sigma)\,p(\mathbf{c},\sigma)$
@@
11:Prior $p(\mathbf{x},\bm{\omega})$, posterior $p(\mathbf{c},\sigma\mid D_{T})$
12:Trained summary and inference network parameters $\widehat{\bm{\theta}},\widehat{\bm{\varphi}}$
@@
15:   Sample inputs and parameters from prior $(\mathbf{x}^{(i)},\bm{\omega}^{(i)})\sim p(\mathbf{x},\bm{\omega})$
@@
17:   Evaluate surrogate $\widetilde{\mathbf{y}}^{(i)}=\widetilde{M}_{\mathbf{c}^{(i)}}(\mathbf{x}^{(i)},\bm{\omega}^{(i)})$
@@
19:   Pass $(\mathbf{x}^{(i)},\widetilde{\mathbf{y}}_{\epsilon}^{(i)})$ through summary network $\mathbf{s}^{(i)}=S_{\bm{\theta}}(\mathbf{x}^{(i)},\widetilde{\mathbf{y}}_{\epsilon}^{(i)})$
20:   Pass $(\bm{\omega}^{(i)},\mathbf{s}^{(i)})$ through inference network $q_{\bm{\varphi}}(\bm{\omega}^{(i)}\mid\mathbf{s}^{(i)})$
21:   Compute loss from Eq.˜2 and update network parameters $\bm{\theta},\bm{\varphi}$